OpenStudy (anonymous):

what do i do it says for the directionsit says: tell whether the two ratios form proportion: it gives me 4/7 and 20/25

OpenStudy (amistre64):

does 4/7 = 20/25 ?

OpenStudy (anonymous):

you know what the directions are right

OpenStudy (amistre64):

a proportion is defined as equal ratios a ratio is defined as a:b, or a/b therefore, if your ratios form a proportion, then they will equal each other.

OpenStudy (amistre64):

i only know what you have typed up as your post :)

OpenStudy (anonymous):

i am confused wait how do i find out if they are equal would i simifly 20/25

OpenStudy (amistre64):

that would be one way you can start it yes

OpenStudy (anonymous):

would that give the awnser

OpenStudy (amistre64):

other ways are if:\[\frac ab = \frac nm\]\[a=b\frac nm\] it would give you something to compare with, and then be able to determine if its true or false

OpenStudy (anonymous):

what does it mean by form a proportions in the question

OpenStudy (amistre64):

to form a proportion, ratios have to be equal. so it is asking if 4/7 is equal to 20/25

OpenStudy (anonymous):

my awnser that i got was